OpenStudy (anonymous):

Can some please help me check over my work: 9: (7c^7d^2)^-2 10. x^5y^-8/x^5y^-6 11. 12a^-3b^9/21a^2b^-5 12. m^-5n^6p^-4/m^-7np^3

OpenStudy (mertsj):

Only if you say what you got can we check the outcome.

OpenStudy (anonymous):

9. ? 10: 1/y^2 11. ? 12. n^5/m^2p^7

OpenStudy (anonymous):

can you help me with 9 & 11

OpenStudy (mertsj):

\[(7c^7d^2)^{-2}=\frac{1}{(7c^7d^2)^2}=\frac{1}{49c ^{14}d ^{4}}\]

OpenStudy (mertsj):

\[\frac{12x ^{-3}b^9}{21a^2b ^{-5}}=\frac{12b^9b^5}{21a^2a^3}=\frac{4b ^{14}}{7a^5}\]

OpenStudy (mertsj):

10 is correct.

OpenStudy (mertsj):

\[\frac{m ^{-5}n^6p ^{-4}}{m ^{-7}np^3}=\frac{m^2n^5}{p^7}\]

OpenStudy (mertsj):

That's number 12. You had it right except for the m^2 is on the top.

OpenStudy (anonymous):

Oh okay, Thank you so muchh!

OpenStudy (mertsj):

yw